Duke Internet Programming Contest

Logistics

November 18, 1993

This document contains 2 pages.^{1}

1 Overview

The Duke Internet Programming Contest is a real-time event. The contest depends on fast, reliable electronic-mail links between your machine and ours. Basically, you run one of three shell scripts on your machine that will send us specially formatted e-mail, which we will then evaluate. It is extremely important that you use the programs we provide, since all of the processing is automated and depends on the special formatting these programs do. Once your request has been processed on our machines, you will receive a reply via e-mail.

If your e-mail link is slow (e.g., you have a UUCP link which only connects once per hour) you will be at a serious disadvantage: you will not receive acknowledgements or clarifications in a timely manner. Also, you will not know what team number you have been assigned until you receive the acknowledgement of your team registration request.

2 Setup and registration

You should begin this process about one hour before the contest begins, but no sooner, please.

Remember that you are only allowed one terminal or workstation per team; essentially, the limit is one keyboard per team. One of the team members should log in as usual. Pick a shell window to use (or, if you are on a terminal, use the login shell) for interacting with the judges. Our examples here assume you are using csh; you may need to adjust if you use another shell. To get started, do:

        add dipc
        register

This will make the contest scripts available to you, and start the registration process. The registration program will ask for your names, questions to determine the class of your team, the name of the site you are from, and your email address. Please be sure that the answers you provide to the register program are correct, especially the e-mail address. If you make a mistake do not confirm the registration. You will receive a message back from the judges confirming your registration. (If you do not receive a message back within 10–15 minutes, see Section 5.) The message will also include your team number. Write down your team number now. All correspondence between us and you is done via the team number, as is the scoring. Follow any other instructions provided by this mail message, or by your local contest administrator. Do not reregister. We highly suggest that you not discard any email you receive from us until the contest is over.

1I have made several changes to the original Duke document for the Athena environment. If you want a copy of the original document, ask me ( Marc Horowitz <marc@mit.edu>).

You might wish to work in a protected directory. To do this, type

   mkdir contest
   fs sa contest -clear $USER all
   cd contest

Now you are ready to begin. The contest will begin precisely at 6pm (1800 hours) Eastern time (2300 GMT, November 18). You should start when your local computer clock reaches that time (try date -u to see GMT time^{2}). Your submissions will be time-stamped from your local computer clock, to help remove network propagation delays from the timing. If your computer clock is off by more than a few minutes from the actual time, please see about having it adjusted by a system adminstrator.

3 To submit a judged run

Once you think you have a problem completed, you need to send the program (source code) to the judges. This is also done through the mail. If your program is called foo.c, you should type

     submit foo.c

This script will ask you for the problem number, your team number, and the language you are using. See the rules for the exact compiler versions we will be using.

You must type yes to confirm your responses (please proofread them first). You will receive mail about the judge’s decision. Do not resubmit the program until you have received the judge’s decision. We will penalize your team for this.

If you have set the environment variable IPCTEAM to your team number, then this program will not ask you to enter the team number.

4 To submit a question for clarification

To ask a question of the judges, type

    clarify

Provide your team number and the problem number. It will put you in your editor (if you have set the environment variable EDITOR) to type your question. Please explain your question clearly. Exit the editor. Confirm the submission, and it will be sent to the judges. The judges will respond, although they may not answer your question. Your question and its answer may be sent to all other teams if it is relevant. You will know whether this is the case.

Clarifications may arrive through the mail at any time. Thus, you should watch your mailbox. If you have set the environment variable IPCTEAM to your team number, then this program will

not ask you to enter the team number.

5 Problems with the contest logistics

Send mail to khera@cs.duke.edu, or call +1-919-660-6581. Do not send mail to contest@cs.duke.edu under any circumstances. If you think your problem is with Athena and not the contest files, send mail to marc@mit.edu.

2I’ve tested this on the vax, dec, sun, and rs/6000. If you have problems, contact me.